OpenStudy (anonymous):

Find the value of the missing numerator and enter it below. 1/x-1/(x+4)= ?/x(x+4)

OpenStudy (campbell_st):

need a common denominator of x(x +4) multiply the 1st term by (x + 4) the 2nd term by x \[\frac{(x +4) \times 1}{x(x+4)} - \frac{x \times (1)}{x(x+4)} = \frac { x - x + 4}{x(x+4)} = \frac{4}{x(x+4)}\]